The length of a rectangle is 20 inches, and the width is four less than the length. Find the area of the rectangle.

Answer by Timnewman(323)

Let L and w be the lenght and the width of the rectangle respectively.
If L=20 ---(1)
w=L-4---(2)
put equ (1) in equ (2)
w=20-4
w=16
Then,Area (A)=L*w
A=20*16
A=320sqrt inch